# Gas-induced perturbations on the gravitational wave in-spiral of live post-Newtonian LISA massive black hole binaries

### v1

@@ 3 Results
Figure 2: Gas torques onto the binary in terms of $\xi$ as a function of the SMA for GW+gas simulation (light blue lines). We show average $\xi$ values for GW+gas (solid blue line; $\bar{\xi}_{\rm GW+gas}\approx-24$) and gas-only (dashed red line; $\bar{\xi}_{\rm gas}\approx-36$) cases, respectively.
Once we subtract both $\dot{a}_{\rm GW}$ and $\dot{a}_{\rm gas}$ from the total SMA $\dot{a}_{\rm GW+gas}$ rate inferred from the full GW+gas simulation, we find a non-negligible residual, which is the GW-gas coupling term introduced in Eq. (3). We propose to express this cross-term simply as
$\dot{a}_{\rm GWgas}\sim\zeta\dot{a}_{\rm GW}\dot{a}_{\rm gas},$ (6)
where $\zeta$ is also a simulation-calibrated parameter and we show its evolution in Fig. 3 as function of SMA. We find the average value of the cross-term to be $\approx-7.9\times 10^{-5}$. The value of $\zeta$ is also decreasing as the binary is shrinking, similar to the evolution of $\xi$ in Fig. 2. The almost constant evolution of $\zeta$, marginally weakened by the in-spiral, may imply that our assumption to express $\dot{a}_{\rm GWgas}\propto\dot{a}_{\rm GW}\dot{a}_{\rm gas}$ is not inaccurate.
Figure 3: GW-gas coupling coefficient as a function of SMA (light blue lines) with its mean value $\bar{\zeta}=-7.9\times 10^{-5}$ (solid blue line).
3.1 Gas-induced dephasing and LISA observability
In this section, we compute the gas-induced dephasing $\delta\phi_{\rm gas}$ in Eq. 4 in the source-frame between $=53.2~{}r_{s}$ and $=30~{}r_{s}$ with the initial eccentricity $=0.29$ by using $\dot{a}_{\rm GW}$ inferred from the simulations and then considering the total SMA evolution rate $\dot{a}_{\rm GW+gas}$ in three different ways.
  • First of all directly from the GW+gas simulations, which leads to a dephasing $\delta\phi_{\rm gas}=-3.21$ radians.
  • The second estimate of the dephasing is performed by neglecting the GW-gas cross term, i.e. assuming $\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}$, where $\dot{a}_{\rm gas}$ is computed from Eq. 2 with the gas-only simulation-calibrated parameters $\xi=-36$ and ${\rm f}_{\rm Edd}=1.2$. This approach gives $\delta\phi_{\rm gas}=-0.21$ radians.
  • Lastly, we include the cross-term using the parameterization in Eq. 6 to have $\dot{a}_{\rm GW+gas}=\dot{a}_{\rm GW}+\dot{a}_{\rm gas}+\zeta\dot{a}_{\rm GW}\dot{a}_{\rm gas}$ with $\zeta=-7.9\times 10^{-5}$. This results in $\delta\phi_{\rm gas}=-3.63$ radians.
The dephasing we obtain neglecting the GW-gas cross term (i.e. assuming $\zeta=0$) is an order of magnitude smaller than the value inferred with the non-zero cross term, so we caution that the GW-gas coupling might have a crucial role in the evolution of MBHBs in gaseous environments. Furthermore, the fact that the $\delta\phi_{\rm gas}$ computed directly from the GW+gas simulation is almost the same as the dephasing computed by parametrizing the gas effect through the (simulation-calibrated) $\xi$ and $\zeta$ parameters, provides a further consistency check of our model. The few percent difference between the two values might originate from the approximation made in assuming $\xi$ and $\zeta$ to be constants with time.
We consider the evolution of our $10^{6}~{}{\rm M}_{\sun}$ total mass eccentric MBHB at redshift $z=1$ for one year of LISA observation. We therefore take an initial SMA of $a_{i}=40r^{z=1}_{s}$ and an initial eccentricity $e=0.2$, as the binary has partially circularized from its initial separation. We find the dephasing for ${\rm f}_{\rm Edd}=1.2$, $\xi=-36$, and $\zeta=-7.9\times 10^{-5}$ between $a_{i}=40r^{z=1}_{s}$ and $a_{f}=30r^{z=1}_{s}$ to be $\delta\phi_{\rm gas}=-1.11$ radians, equivalent to losing $\sim 0.2$ GW cycles over the expected $\sim 850$ cycles from a vacuum eccentric binary evolving just under GWs. Since LISA should observe this event with SNR1300similar-toabsent1300$\sim$ 1300∼ 1300 (Garg et al., 2024d) and we typically need a given dephasing to be higher than 8$\sim$ 8/∼ 8 /SNR (Derdzinski et al., 2021; Garg et al., 2022) to be detectable, we should be able to observe the computed $\delta\phi_{\rm gas}$.
3.1.1 Analytical dephasing estimates
Since we would ultimately like to obtain a parametrized model for the GW+gas inspiral phase, we now use the orbital-averaged GW-driven inspiral rate at $2.5$PN order from Peters (1964), i.e. $\dot{a}^{\rm Peters}_{\rm GW}$, for the GW only driven inspiral rate. Moreover, we also analytically construct the total SMA rate using the simulation-calibrated parameters $\xi$, ${\rm f}_{\rm Edd}$, and $\zeta$ to have
$\dot{a}^{\rm analytical}_{\rm GW+gas}=\dot{a}^{\rm Peters}_{\rm GW}+\dot{a}_{\rm gas}+\zeta\dot{a}^{\rm Peters}_{\rm GW}\dot{a}_{\rm gas}.$ (7)
If we repeat the dephasing computation done in § 3.1, then we get $\delta\phi_{\rm gas}=-3.75$ radians. This is a few percent higher than the value calculated using only the simulations data. One possible reason is the inclusion of 1PN and 2PN terms in the simulations which, despite being conservative terms, lead to the precession of the MBHB eccentric orbit effectively altering its in-spiral rate. Another source of uncertainty in the dephasing estimate might reside in numerical errors. We have however compared our GW-only simulation with a numerical integration of the full (i.e. up to 2.5 order) PN evolution of the binary using an 8th order Runge-Kutta integrator finding negligible differences within the separations range we explored.
Even though we have only performed simulations for one set of binary-disk parameters, we can compute the expected dephasing for different values of the $\xi$ and $\zeta$ parameters, which reflect the use of different system parameters in the simulations, through the analytical prescription in Eq. 7. In Fig. 4, we show $\delta\phi_{\rm gas}$ for equal-mass circular MBHBs accreting at ${\rm f}_{\rm Edd}=1$ for a one-year LISA observation until respective ISCO with varying $\xi\in[-1,-100]$, $\zeta\in[0,2\times 10^{-4}]$, $M\in\{10^{5},10^{6}\}~{}{\rm M}_{\sun}$, and $z\in\{1,2\}$ to bracket reasonable parameters space. If we do not consider the GW-gas cross-term then the dephasing cannot be detectable by LISA for $M=10^{6}~{}{\rm M}_{\sun}$ MBHB at both redshifts for $|\xi|\lesssim 1$. The stronger dephasing with non-zero $\zeta$ again signifies the importance of considering the GW-gas cross-term in having observable gas effects on the GW waveform. Moreover, similar $\delta\phi_{\rm gas}$ for different combinations of mass and redshift is possibly due to the cross-term dependence on $\dot{a}^{\rm Peters}_{\rm GW}$.
Figure 4: The gas-induced dephasing in the GW waveform, $\delta\phi_{\rm gas}$ in radians, over a one-year LISA observation until ISCO as a function of the simulation-calibrated fudge factors $\xi$ (represents gas torques) and $\zeta$ (represents GW-gas cross-term) for equal-mass MBHBs accreting at ${\rm f}_{\rm Edd}=1$. In all four panels, we vary $\xi$ from $-1$ to $-100$ and $\zeta$ from 00 to $2\times 10^{-4}$. In the top two panels, we consider systems at redshift $z=2$ and in the bottom two panels we assume $z=1$. In the left two panels, we study a MBHB with source-frame total mass $M=10^{5}~{}{\rm M}_{\sun}$ and in the right panels we have $M=10^{6}~{}{\rm M}_{\sun}$. We also draw cyan lines to indicate the regions of non-observable dephasing by LISA, i.e. $\delta\phi_{\rm gas}<8/$SNR.
A perturbative correction to $\dot{a}_{\rm GW}$ with a functional form $\propto a^{\rm ng}\dot{a}_{\rm GW}$ manifests at $-n_{g}$ relative order in the GW phase (Garg et al., 2024a). We have $n_{g}=4$ and $n_{g}=1$ for $\dot{a}_{\rm gas}$ and $\dot{a}_{\rm GWgas}$, respectively. Hence, an instantaneous frequency-depending dephasing enters the overall binary phase at relative $-4$PN order due to evolution just under gas and at $-1$PN relative order due to a GW-gas coupling.
Lastly, if we do not introduce a GW-gas coupling and try to explain the residual $(\dot{a}_{\rm GW+gas}-\dot{a}_{\rm GW})$ by using the model for $\dot{a}_{\rm gas}$ in Eq. 2, then the mean value of the parameter $\xi$ is $\sim-360$. This implies that even a dephasing at $-4$PN relative order should be observable for our studied MBHB with or without concurrent eccentricity measurement (Garg et al., 2024a) in a realistic Bayesian data analysis setup, where all other binary parameters are also varying and a complete LISA response is considered. However, the GW-gas coupling is even more stronger than this, since it enters the GW phase at $-1$PN relative order, therefore causing a larger dephasing.
@@ 4 Discussion and Conclusion
We studied the interaction of an equal mass MBHB with its surrounding CBD during the late in-spiral stage, with and without concurrent GW emission, using 3D hydrodynamical simulations with a live binary. This approach provided us with the first direct measurement of how surrounding gas torques the binary when its in-spiral is already governed by GW emission, by means of the estimate of the $\xi$ parameter in Eq. (1). We find that $\xi$ is $\mathcal{O}(10)$ stronger than the previously explored scale-free/sub-pc regime. However, we caution that the comparison may not be fair as previous simulations carried out for the larger separation regime are predominantly 2D, assume a fixed binary orbit modelled under Newtonian dynamics, and compute the effect of the energy and angular momentum loss by GW radiation in post-processing (Tang et al., 2018). In particular, Duffell et al. (2024) showed that 3D calculations give different magnitude torques compared to 2D and Franchini et al. (2023) argued that fixing the binary orbit leads again to a different gravitational torque. Moreover, we find that, consistently with our expectations, $\xi$ is slightly weaker for GW+gas simulations ($\xi_{\rm GW+gas}\sim-24$) with respect to gas-only study ($\xi_{\rm GW+gas}\sim-36$), as shown in Fig. 2. This is expected since the binary is decoupling from the gas in the GW+gas simulation and therefore the effect of gas weakens with time.
Most importantly, we show that our results can be understood by introducing a novel coupling term between the energy loss by GWs and that due to gas in the orbital in-spiral rate, which we model through the product between the rate of change of the SMA due to GWs only and that due to gas only. The amplitude of this coupling term is pinpointed by a newly introduced simulation-calibrated parameter $\zeta$, whose evolution is shown in Fig. 3 with an average value $\sim-7.9\times 10^{-5}$. This implies richer gas features in the GW waveform than previously expected. While gas torques still induce a correction to the GW phase at the usual $-4$PN relative order (Garg et al., 2024a), our modeling of GW-gas cross-term enters the phase at $-1$PN relative order, leading to stronger phase shift by $\mathcal{O}(10)$ and higher likelihood of detection of gas perturbations with LISA. The $-1$PN relative order term may also allow us to break the degeneracy with concurrent eccentricity measurement using LISA (Garg et al., 2024a). Moreover, having a GW-gas coupling may change the expectation how the binary parameters in the LISA band are correlated with their values at sub-pc scales (Garg et al., 2024c). Lastly, ignoring gas effects while modeling MBHBs may lead to stronger apparent violations of general relativity than previously anticipated (Garg et al., 2024b) due to the new $-1$PN relative order term.
In this work, we studied one set of binary-disk parameters while in the future we plan on exploring the parameter space to understand how both $\xi$ and $\zeta$ change. This exercise may also allow us to analytically model both simulation-calibrated parameters $\xi$ and $\zeta$. For instance, if the binary is embedded in a relatively thick disk ($H/R=0.1$), we can expect to also have a significant contribution from mini-disks (i.e. disks around each binary component) that will introduce their own likely positive torque measurement (D’Orazio & Duffell, 2021; Franchini et al., 2022; Siwek et al., 2023; Tiede & D’Orazio, 2023) and probably a different sign for $\zeta$. Moreover, different numerical studies may also allow to check the robustness of our phenomenological model.
Lastly, we also find that, in the gas-only simulation, the eccentricity is linearly decreasing with the SMA. This needs to be further investigated in both sub-pc and near-merger regimes, and we plan to do so in the future.
@@
It is to be noted that we employ PN corrections as an approximation since we do not run general relativistic magnetohydrodynamical (GRMHD) simulations. Current GRMHD simulations only study the binary evolution just a few days before merger (Gutiérrez et al., 2022; Avara et al., 2023) to integrate only a few orbits because of the prohibitive computational cost. Therefore, since the majority of the gas-driven effects on the binary in-spiral occurs at separations $a\gtrsim 30~{}r_{s}$, our approach is currently the best available method to investigate the orbital dephasing due to the presence of the gas and where PN corrections are may be adequate.
@@
A possible caveat in our work is that our binary is moderately eccentric ($\sim 0.3$) just a few years before the merger, which will require the eccentricity to be extremely high when GWs take over at milli-pc scales. However, since $e\sim 0.3$ arises naturally from our initial condition requiring a steady-state disk before setting the physical scale of $a\sim 53r_{s}$, the only truly realistic way to initialize the system is to evolve the binary starting from a much wider separation. This, however, would increase the computational cost dramatically. We plan to investigate alternative procedures in the setup of the simulations in order to reduce the computational cost and mitigate this issue in the future.
@@
Acknowledgements
MG acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092. The authors also acknowledge use of the NumPy (Harris et al., 2020) and Matplotlib (Hunter, 2007).

### v3

@@ 3 Results
Figure 2: Gas torques onto the binary in terms of $\xi$ as a function of the SMA for gas+PN2.5 simulation (light blue lines). We show average $\xi$ values for gas+PN2.5 (solid blue line; $\bar{\xi}_{\rm{gas+PN2.5}}\approx-19.5$) and gas+PN2 (dashed red line; $\bar{\xi}_{\rm gas+PN2}\approx-24.5$) cases, respectively. Note that we have not done any smoothing in plotting $\xi$ for the gas+PN2.5 run but only interpolation between snapshots.
We also note that the high-frequency or sub-orbital fluctuations in Fig. 2 around the mean value may be measurable on their own via GWs, as suggested by analytical studies (Zwick et al., 2022, 2024, 2025). However, recently, Copparoni et al. (2025) demonstrated, using realistic LISA data analysis, that these moderate fluctuations we find are not measurable in GWs, albeit for a much smaller mass ratio BHB compared to the one explored in our work.
We show torque density maps averages over ten snapshots for the gas+PN2.5 (left panel) and gas+PN2 (middle panel) simulations together with their difference (right panel) in Fig. 3. Similar to previous HD studies (see, e.g. Tiede et al., 2020; Franchini et al., 2023), we rotate gas particles about the binary’s center-of-mass such that the axis connecting the two BHs aligns in the two simulations to do a meaningful comparison between their gas distribution. We chose to average over ten orbits, as it is small enough that the binary is not significantly shrinking and large enough to remove instantaneous features. The first two panels show the contribution from both CBD and mini-disks to the gravitational torque. The gas in front of the BH tends to drag it (pink color) along while the gas behind it decelerates it (green color). The right panel clearly shows that the gravitational torque is larger in the gas+PN2 simulation as there are more regions where the torque is negative. This is again consistent with the estimate of the $\xi$ parameter in Fig. 2.
Figure 3: 2D projected gravitational torque ($T_{\rm grav}$) distribution between 5a-5a to $5a$ in both axes averaged over $100$ snapshots between $100^{\rm th}$ and $110^{\rm th}$ orbits. We show results from the gas+PN2.5 run (left panel), the gas+PN2 run (middle panel), and their difference (right panel). The third panel clearly shows that the gas+PN2.5 run has slightly more positive torque than the gas+PN2 simulation.
3.1 Gas-induced dephasings and LISA observability
In this section, we consider different dephasings introduced at the end of § 2.1. We enumerate source-frame gas-induced dephasings that corresponds to elapsed physical time between $54.5~r_{s}$ and $48.9~r_{s}$ for the gas+PN2.5 simulation over 278 orbits or $\sim 1745.8$ rad orbital phase in Table 1. We have compared our PN2.5 simulation with a numerical integration of the full (i.e. up to 2.5 order) PN evolution of the binary using an 8th order Runge-Kutta integrator finding mean relative error of $\lesssim 0.1\%$ within the separations range we explored. This illustrates that our orbital parameters are well measured in different simulations.
  $\rm\bf Dephasing$}   $\bf Value[rad]$}
  $\delta\phi_{\rm orb}^{\rm(GW)}$   0.007-0.007
  $\delta\phi_{\rm ecc}^{\rm(GW)}$   0.012-0.012
  $\delta\phi_{\rm orb}^{\rm(NoGW)}$   0.014-0.014
  $\delta\phi_{\rm ecc}^{\rm(NoGW)}$   0.010-0.010
  $\delta\phi_{\rm GW}$   0.014-0.014
Table 1: Dephasings measured from our simulations: gas-induced orbital $\delta\phi_{\rm orb}^{\rm(GW)}$ and precession $\delta\phi_{\rm ecc}^{\rm(GW)}$ dephasings between gas+PN2.5 vs PN2.5 simulations and same dephasings but between gas+PN2 and PN2 runs ($\delta\phi_{\rm orb}^{\rm(NoGW)},\delta\phi_{\rm ecc}^{\rm(NoGW)}$). We also infer gas-induced GW dephasing by doubling $\delta\phi_{\rm orb}^{\rm(GW)}$.
The comparison between the orbital phase evolution in the gas+PN2.5 and PN2.5 simulations, $\delta\phi_{\rm orb}^{\rm(GW)}$, allows us to quantify the overall contribution of the different perturbations induced by the gas disk in the GW-driven binary evolution, i.e. its effect on $\dot{a}$, $\dot{e}$, and the precession rate. We can directly measure the dephasing due to the different precession rates $\delta\phi_{\rm ecc}^{\rm(GW)}$ directly from the simulations. However this is still not completely independent from other changes in the orbital parameters. Moreover, GWs radiation further complicates the picture as it changes the gas morphology, which can further contribute to $\delta\phi_{\rm orb}^{\rm(GW)}$ in a non-linear manner.
Comparing the gas+PN2 and the PN2 runs can alleviate some of these issues as these runs both neglect the GW-driven fast inspiral, allowing us to measure the dephasing due to disk-induced binary precession $\delta\phi_{\rm ecc}^{\rm(NoGW)}$ in a system where the semi-major axis and eccentricity of the binary do not significantly evolve with time.
The results presented in Table 1 show that the orbital dephasing measured in the gas+PN2.5 simulation (first row) is smaller than the same dephasing inferred in absence of GW emission (third row). This is consistent with the effect of the gas becoming weaker as the binary inspirals. The estimate of dephasing due to the binary precession in the second row is affected by the change in binary eccentricity induced by GW emission and is therefore not directly comparable to the term in the first row. The precession-induced dephasing calculated from the NoGW simulations (fourth row) is more similar to the term in the first row. This seems to indicate that the main contributor to the binary orbital dephasing is the disk-induced precession of the binary eccentricity vector. However we restrain to draw such a strong conclusion as longer simulations together with a larger exploration of the parameter space is needed to understand the role that non-linear effects play in the ultimate orbital dephasing in GWs. Furthermore, we note that all the dephasing inferred from the simulations are subjected to the particles noise and to the precise alignment of the orbital phase and argument of periapsis at the end of the simulations.
If we compute the gas-induced GW dephasing analytically using only the gas torques by linearly adding SMA rates due to GWs and gas (Garg et al., 2022; Dittmann et al., 2023; Duffell et al., 2024) then we get $\delta\phi_{\rm GW}^{(\xi)}=-0.095$ rad, which is a factor of $7$ higher than our direct estimate of $\delta\phi_{\rm GW}=-0.014$ rad. This difference could be due to i) the inadequacy of the analytical prescription, ii) the approximation made by simply linearly adding two terms (i.e. GW-driven and gas-driven SMA change) that mutually affect each other and iii) different numerical approaches. Follow-up simulations might shed more light onto the nature of this discrepancy. We note that our estimate of precession-induced dephasing of $\delta\phi_{\rm ecc}^{\rm(GW)}=-0.012$ rad over 278 orbits is in magnitude comparable to the analytical prescription by Tiede et al. (2024) that yields dephasing of $\sim 0.010$ rad. This indicates that the choice of numerical method has a limited effect on the precession-induced dephasing.
[…]
@@ 4 Discussion and Conclusion
We studied the interaction of an equal mass MBHB with its surrounding geometrically thin ($H/R=0.03$) CBD during the late in-spiral stage, with and without concurrent GW emission, using 3D hydrodynamical simulations with a live binary. This approach provided us with the first direct measurement of how surrounding gas torques the binary when its in-spiral is already governed by GW emission, by means of the estimate of the $\xi$ parameter in Eq. (1). We find that $\xi$ is $\mathcal{O}(10)$ stronger than in some of the previously explored scale-free/sub-pc regime. However, we caution that the comparison may not be fair as previous simulations carried out for the larger separation regime are predominantly 2D, assume a fixed binary orbit modeled under Newtonian dynamics, explore larger values of the viscosity parameter $\nu$ (Dittmann & Ryan, 2022), and compute the effect of the energy and angular momentum loss by GW radiation in post-processing (Tang et al., 2018).
Note that if we compared our value of $\xi$ in the gas+PN2 simulation with the results presented in Tiede et al. (2025) with the same value of viscosity that we have and the highest resolution they explored, we find the difference to be a factor of two. We also find that the measured time-averaged torque $\xi$ becomes weaker with higher resolutions (see Fig. A1), which is the opposite trend to the one found by Tiede et al. (2025). The different behavior might be due to our 3D live-orbit PN treatment with respect to their 2D fixed-orbit Newtonian simulations and to the different Mach number, i.e. $\mathcal{M}\approx 33$, employed in our simulations rather than their $\mathcal{M}=40$. In particular, Duffell et al. (2024) showed that 3D calculations give different magnitude torques compared to 2D and Franchini et al. (2023) argued that fixing the binary orbit leads again to a different gravitational torque. Moreover, we find that, consistently with our expectations, $\xi$ is slightly weaker for gas+PN2.5 simulations ($\bar{\xi}_{\rm{gas+PN2.5}}\sim-19.5$) with respect to gas+PN2 study ($\bar{\xi}_{\rm{gas+PN2}}\sim-24.4$), as shown in Fig. 2. This is expected since the binary is decoupling from the gas in the gas+PN2.5 simulation and therefore the effect of gas weakens with time.
@@
It is to be noted that we employ PN corrections as an approximation since we do not run general relativistic magnetohydrodynamical (GRMHD) simulations. Current GRMHD simulations only study the binary evolution just a few days before merger (Gutiérrez et al., 2022; Avara et al., 2024) to integrate only a few orbits because of the prohibitive computational cost. Therefore, since the majority of the gas-driven effects on the binary in-spiral occurs at separations $a\gtrsim 48.9~r_{s}$, our approach is currently the best available method to investigate the orbital dephasing due to the presence of the gas and where PN corrections may be adequate.
A caveat in this work is that we assume that the gas morphology will look the same at our initial separation for all our gas+PN2.5 and gas+PN2 simulations. In principle, one would need to start the binary at sufficiently large separations, approaching parsecs, that the GW radiation is completely negligible, and quantify the difference in the morphology of the gas distribution while the binary shrinks as the GW radiation gradually ensues relative to a case in which it is neglected. This is currently not possible due to the prohibitive computational costs. If anything, by starting the binary with the same initial condition in the gas+PN2.5 and gas+PN2 simulation at a separation at which GW radiation is already taking place, we are erring on the side of underestimating the back-reaction of the gas to the GW emission, which translates into a conservative estimate of the cross-term.
@@
Another possible caveat in our work is that our binary is moderately eccentric ($\sim 0.3$) just a few years before the merger, which will require the eccentricity to be extremely high when GWs take over at milli-pc scales. However, since $e\sim 0.3$ arises naturally from our initial condition requiring a steady-state disk before setting the physical scale of $a\sim 55r_{s}$, the only truly realistic way to initialize the system is to evolve the binary starting from a much wider separation. This, however, would increase the computational cost dramatically. We plan to investigate alternative procedures in the setup of the simulations in order to reduce the computational cost and mitigate this issue in the future.
@@
Acknowledgments
MG acknowledge support from the Swiss National Science Foundation (SNSF) under the grant 200020_192092 and “GW-learn” grant agreement CRSII5 213497. AF acknowledges support provided by the “GW-learn” grant agreement CRSII5 213497 and the Tomalla Foundation. AL acknowledges support by PRIN MUR “2022935STW”. We thank the anonymous referee for helpful comments that improved this work. We futher thank Rohit Chandramouli, Andrea Derdzinski, Alexander Dittmann, Callum Fairbairn, Zoltan Haiman, Laura Sberna, Connar Rowan, Christopher Tiede, and Lorenz Zwick for useful discussions. The authors also acknowledge use of the NumPy (Harris et al., 2020) and Matplotlib (Hunter, 2007).
Appendix A Resolution study
Figure A1: Average torque value (blue cross) expressed in terms of $\bar{\xi}$ over initial 100 orbits of gas+PN2.5 simulations for three different resolutions: LR with $\Delta x[3a]={1.31~r_{s}}$, MR with $\Delta x[3a]={1.07~r_{s}}$, and HR with $\Delta x[3a]={0.66~r_{s}}$.
In Fig. A1, we show the torque values $\bar{\xi}$ time-averaged over 100 initial binary orbits, or equivalently in the SMA range between $54.5~r_{s}$ and $52.6~r_{s}$, for three different resolutions of the gaseous disk. We quantify the resolution by measuring an equivalent inter-particle spacing $\Delta x[3a]$ evaluated at $R=3a$. We name the three simulations as: low-resolution (LR) with $\Delta x[3a]={1.31~r_{s}}$, mid-resolution (MR) with $\Delta x[3a]={1.07~r_{s}}$, and high-resolution (HR) with $\Delta x[3a]={0.66~r_{s}}$. Fig. A1 shows the value of $\bar{\xi}$ for the three resolutions. Since the values measured from our MR and HR runs are very similar, within $0.4\%$, we can conclude that the MR run is sufficiently converged and we therefore further evolve it until $278$ orbits to measure the dephasing induced by the interaction of the binary with the disk. We run the gas+PN2 run at the MR resolution of $\Delta x[3a]={1.07~r_{s}}$ for the same elapsed physical time in order to measure the differences between the two simulations (see Section § 3). Since the gas+PN2 MR setup has no fast inspiral due to GWs, we can reasonably assume it is also converged.